OpenStudy (anonymous):

A new type of pump can drain a certain pool in 10 hours. An older pump can drain the pool in 15 hours. How long will it take both pumps working together to drain the pool?

OpenStudy (stefrheart):

First you set up the equation as \[\frac{ 1 }{ 10 } + \frac{ 1 }{ 15 } \] =T

OpenStudy (stefrheart):

Then you add the two fractions i would make the denominator 30 since that would be the easiest \[\frac{ 3 }{ 10 } + \frac{ 2 }{ 30 }\]

OpenStudy (stefrheart):

After adding that it should look like this \[\frac{ 5 }{ 30 }=\frac{ 1 }{ t }\] then flip both so it becomes \[\frac{ 30 }{ 5 } = t\] then solve and you should get your answer also sorry i made a mistake earlier when i put it equaled to t on the first one, it should have been 1/t

OpenStudy (anonymous):

Its all good, really appreciate it man

OpenStudy (stefrheart):

Welcome! Glad to help :)